# No-go theorem for probabilistic one-way secret-key distillation

### v1

@@ Abstract
The probabilistic one-way distillable secret key is equal to the largest expected rate at which perfect secret key bits can be probabilistically distilled from a bipartite state by means of local operations and one-way classical communication. Here we define the set of super two-extendible states and prove that an arbitrary state in this set cannot be used for probabilistic one-way secret-key distillation. This broad class of states includes both erased states and all full-rank states. Comparing the probabilistic one-way distillable secret key with the more commonly studied approximate one-way distillable secret key, our results demonstrate an extreme gap between them for many states of interest, with the approximate one-way distillable secret key being much larger. Our findings naturally extend to probabilistic one-way entanglement distillation, with similar conclusions.
preprint:
LABEL:FirstPage1 LABEL:LastPage#110
Introduction—Quantum key distribution has emerged as one of the most promising applications of a quantum network, as it facilitates unconditionally secure communication between distant parties [1, 2, 3]. It allows the transmission of private data between multiple parties, such that the security is ensured by the laws of quantum mechanics [4, 5], instead of relying on computational assumptions about the eavesdropper [6]. The rapid development of quantum network technologies demands a strong understanding of our ability to distribute a secret key over a quantum network equipped with some available resources.
@@
In this paper, we analyze the probabilistic secret-key distillation approach in which a perfect secret key is distilled from an initial bipartite state, albeit probabilistically. In particular, we study the one-way distillable secret key of a bipartite state in the probabilistic setting, which is roughly defined as the maximum achievable expected rate of distilling secret key bits from an arbitrarily large number of copies of the state using one-way LOCC channels.
@@
One of the main tools that we use to establish our results is the resource theory of unextendibility [20, 21] and its state-dependent variation [22]. This resource theory was developed in [20, 21] as a relaxation of the resource theory of entanglement, in which one-way LOCC channels are allowed for free. An important quantity that we employ is the min-unextendible entanglement of a bipartite state [22]. Our work also provides significant improvements on existing bounds [22, Sec. V.D] regarding the overhead of probabilistic one-way secret key distillation.
One-way secret-key distillation—Let us begin by considering the task of secret-key distillation. The objective of such a protocol $\mathcal{L}_{AB\to A^{\prime}B^{\prime}}$ is to transform a bipartite state $\rho_{AB}$, purified by $\psi_{ABE}$, into a tripartite key state as follows:
@@
where $\omega_{A_{1}B_{1}}$ is an arbitrary quantum state and $(U_{i})_{i=0}^{k-1}$ is a tuple of unitary operators. The systems $A_{0}B_{0}$ are the key systems, and $A_{1}B_{1}$ are the shield systems. The bipartite private state defined in (2) can be used to distill at least $\log_{2}k$ secret key bits. We can hence reframe the task of secret key distillation into the distillation of bipartite private states. From here on we simplify the labeling for the key systems and shield systems: when referring to a private state $\gamma^{k}_{A_{0}A_{1}B_{0}B_{1}}$, we group the systems held by Alice into one system label $A\coloneqq A_{0}A_{1}$ and all the systems held by Bob into one system label $B\coloneqq B_{0}B_{1}$.
@@
where we have used the shorthand $[i]\coloneqq|i\rangle\!\langle i|$. In (3), $\gamma^{k}_{A^{\prime}B^{\prime}}$ is a bipartite private state with at least $\log_{2}k$ bits of secrecy, system $X$ is a classical flag indicating the success or failure of the protocol, and $\sigma_{A^{\prime}B^{\prime}}$ is an arbitrary quantum state generated when the protocol fails to distill a private state.
In any probabilistic secret-key distillation protocol, both parties must have access to the flag $X$ in order to use the distilled key for private communication. Suppose a probabilistic secret-key distillation protocol fails to establish a secret key. In that case, both parties involved in the distillation process can discard their systems and repeat the protocol with another instance of the resource state. If one-way LOCC channels are available for free, it suffices to demand that Alice receives the flag $X$ since she can send the flag to Bob using the freely available forward classical channel.
@@
In (5) we require that the state generated upon failure of the protocol is a maximally mixed state. This additional constraint on secret-key distillation protocols does not affect the maximum expected number of secret key bits that can be distilled from a bipartite state using one-way LOCC channels as we show in Appendix A.
It is worthwhile to note that a maximally entangled state of Schmidt rank $k$ is a private state holding $\log_{2}k$ secret key bits [8, 9]. As such, any entanglement distillation protocol can be transformed into a secret-key distillation protocol without affecting the rate of distillation. Hence, the one-way distillable secret key of a quantum state is not less than the one-way distillable entanglement of the state in both the probabilistic and approximate settings.
A simpler way to analyze probabilistic distillation of private states under the action of one-way LOCC channels is by considering the erasure symbol. If Alice finds the flag $X_{A}$ in the state $[0]_{X_{A}}$, she can erase her state, and she can instruct Bob to erase his state as well. We call the resulting state the doubly erased private state, which has the following form:
$\eta^{p,k}_{AB}\coloneqq p~{}\gamma^{k}_{AB}+(1-p)[e]_{A}\otimes[e]_{B},$ (6)
where $[e]$ is the erasure symbol that is orthogonal to every state in the span of $\{|i\rangle\!\langle j|\}_{i,j=0}^{d-1}$ and $d$ is the dimension of the underlying system $A$ or $B$. In case of a joint system $S$, the state $[e]_{S}$ represents a state that is orthogonal to every state in the Hilbert space of the system $S$. Both Alice and Bob can retrieve the flag by performing the POVM $\{\Pi,[e]\}$ on their respective systems, where $\Pi\coloneqq\sum_{i=0}^{d-1}[i]$. They can further replace their state with a maximally mixed state upon measuring the erasure symbol, hence retrieving the distilled state in (5). That is,
$\eta^{p,k}_{AB}\xleftrightarrow{\operatorname{LO}}p[1]_{X_{A}}\otimes\gamma^{k}_{A^{\prime}B^{\prime}}+(1-p)[0]_{X_{A}}\otimes\frac{I_{A^{\prime}B^{\prime}}}{d_{A^{\prime}}d_{B^{\prime}}}.$ (7)
@@
Min-unextendible entanglement—The min-unextendible entanglement of a bipartite state was defined in [22, Sec. 4.2] as a monotone for the state-dependent resource theory of unextendibility. We briefly mention the relevant properties of this quantity here, and we refer the reader to [22] for a detailed presentation of the resource theory of unextendibility.
The min-unextendible entanglement is defined with respect to the min-relative entropy [23, Def. 2] as follows:
@@
$\mathcal{F}\!\left(\rho_{AB}\right)\coloneqq\left\{\begin{array}[]{c}\operatorname{Tr}_{B}\!\left[\rho_{ABB^{\prime}}\right]:\par\rho_{AB}=\operatorname{Tr}_{B^{\prime}}\!\left[\rho_{ABB^{\prime}}\right],\\ \rho_{ABB^{\prime}}\in\mathcal{S}\!\left(ABB^{\prime}\right)\end{array}\right\},$ (9)
with $\mathcal{S}\!\left(ABB^{\prime}\right)$ the set of all states of the joint system $ABB^{\prime}$.
@@
$E^{u}_{\min}\!\left(\eta^{p,k}_{AB}\right)=E^{u}_{\min}\!\left(p[1]_{X_{A}}\otimes\gamma^{k}_{AB}+(1-p)[0]_{X_{A}}\otimes\frac{I_{AB}}{d_{A}d_{B}}\right).$ (10)
Limitations on probabilistic one-way distillable secret key—The monotonicity of the min-unextendible entanglement of a bipartite state under the action of one-way LOCC channels implies that the min-unextendible entanglement of the target state in a probabilistic one-way secret-key distillation protocol does not exceed the min-unextendible entanglement of the source state. We first present a lower bound on the min-unextendible entanglement of the doubly erased private state, which is the target state of a probabilistic one-way secret-key distillation protocol.
@@
The min-unextendible entanglement of the doubly erased private state $\eta^{p,k}_{AB}$ is strictly positive for all $p\in(0,1]$ and every integer $k\geq 2$, and it is only equal to zero at $p=0$ because the resulting state is a product state ([22, Proposition 3]). As a consequence of the one-way LOCC monotonicity of the min-unextendible entanglement, a state whose min-unextendible entanglement is equal to zero cannot be used to distill any number of secret key bits with a non-zero probability using one-way LOCC channels. This is formally stated and proved in what follows.
Let us first analyze the set of states whose min-unextedible entanglement is equal to zero. If there exists a state $\sigma_{AB}\in\mathcal{F}\!\left(\rho_{AB}\right)$ such that $\rho_{AB}=\sigma_{AB}$, then $\rho_{AB}$ is a two-extendible state [24, 25, 26], and its min-unextendible entanglement is equal to zero. A more general set of states for which the min-unextendible entanglement is equal to zero, which we call super two-extendible states, can be defined as follows:
@@
The set of super two-extendible states is convex and open (see Appendix D).
@@
While the approximate one-way distillable secret key of a two-extendible state, also known as an anti-degradable state [27], is equal to zero [28, Thm. 15.43], the same is not true for a general super two-extendible state, as we shall see later in this paper. However, a super two-extendible state cannot be used for probabilistic one-way secret key distillation since its min-unextendible entanglement is equal to zero. Combining this fact with the additive property of the min-unextendible entanglement, we arrive at our main no-go theorem stated as Theorem 1, which identifies a broad set of states for which the probabilistic one-way distillable secret key is equal to zero.
@@
Erased private state—Let us consider the following state, which is established if Bob’s share of a bipartite private state $\gamma^{k}_{AB}$ gets erased with probability $1-p$:
$\widetilde{\eta}^{p,k}_{AB}\coloneqq p~{}\gamma^{k}_{AB}+(1-p)\operatorname{Tr}_{B}\!\left[\gamma^{k}_{AB}\right]\otimes[e]_{B},$ (13)
@@
$\widetilde{\Phi}^{p,d}_{AB}\coloneqq p~{}\Phi^{d}_{AB}+(1-p)\frac{I_{A}}{d_{A}}\otimes[e]_{B},$ (14)
@@
The class of full-rank states also is in the set of super two-extendible states, which leads to Corollary 2 below.
@@
The inability to probabilistically distill resource from full-rank states has been observed in general quantum resource theories [29, 30, 31]. Corollary 2 extends this result to probabilistic secret-key distillation under one-way LOCC channels. This result also implies that the overhead of probabilistic secret-key distillation considered in [22, Sec. V-B] is infinite for quantum states with full-rank density matrices, thus providing a significant strengthening of the bounds from [22, Sec. V-B].
Approximate vs. probabilistic secret key distillation—The probabilistic one-way distillable secret key is demonstrably different from the more commonly studied approximate one-way distillable secret key of bipartite states. The approximate one-way distillable secret key is not less than the approximate one-way distillable entanglement of a state, which in turn is bounded from below by the coherent information of the state [11, Thm. 10]. For $p>\frac{1}{2}$, the coherent information of an erased state $\widetilde{\Phi}^{p,d}_{AB}$ is strictly non-negative. Similarly, the coherent information of some isotropic [32] and Werner states [33], which are full-rank states, is non-zero (see Appendices I and J). However, Corollaries 1 and 2 state that the probabilistic one-way distillable secret key of these states is equal to zero, demonstrating an extreme gap between the probabilistic and approximate one-way distillable secret key.
Conclusion—We found a convex and open set of states, that we call super two-extendible states, from which it is impossible to probabilistically distill secret key bits with any non-zero probability. Our main result is a no-go theorem stating that the probabilistic one-way distillable secret key of a super two-extendible state, and consequently its probabilistic one-way distillable entanglement, is equal to zero.
We demonstrated an extreme gap between the probabilistic and approximate one-way distillable secret key for some states. Considering the intermediate regime between probabilistic and approximate distillation [30, 34, 35], where the goal is to distill high-fidelity maximal-resource states with a non-zero probability, should interpolate between the two extreme cases.
Acknowledgements.

### v2

# No-go theorem for heralded exact one-way key distillation

@@ Abstract
The heralded exact one-way distillable secret key is equal to the largest expected rate at which perfect secret key bits can be probabilistically distilled from a bipartite state by means of local operations and one-way classical communication. Here we define the set of super two-extendible states and prove that an arbitrary state in this set cannot be used for heralded exact one-way secret-key distillation. This broad class of states includes both erased states and all full-rank states. Comparing the heralded exact one-way distillable secret key with the more commonly studied approximate one-way distillable secret key, our results demonstrate an extreme gap between them for many states of interest, with the approximate one-way distillable secret key being much larger. Our findings naturally extend to heralded exact one-way entanglement distillation, with similar conclusions.
1 Introduction
Quantum key distribution has emerged as one of the most promising applications of a quantum network, as it facilitates unconditionally secure communication between distant parties [1, 2, 3]. It allows the transmission of private data between multiple parties, such that the security is ensured by the laws of quantum mechanics [4, 5, 6], instead of relying on computational assumptions about the eavesdropper [7]. The rapid development of quantum network technologies demands a strong understanding of our ability to distribute a secret key over a quantum network equipped with some available resources.
@@ 1 Introduction
In this paper, we analyze the heralded exact one-way secret-key distillation approach in which a perfect secret key is distilled from an initial bipartite state, albeit probabilistically. We also refer to this setting as probabilistic secret-key distillation with zero error, and we adopt the latter terminology henceforth. In particular, we study the one-way distillable secret key of a bipartite state in the probabilistic setting, which is roughly defined as the maximum achievable expected rate of distilling secret key bits from an arbitrarily large number of copies of the state using one-way LOCC channels.
Sometimes the term probabilistic resource distillation is used for probabilistic approximate distillation of a resource in the asymptotic setting, for example in [21, 22]. In that setting, a sequence of probabilistic protocols is considered a valid probabilistic resource distillation protocol if the sequence of distilled states, when the protocol is successful, converges to the maximal-resource state asymptotically. We emphasize that we use the term probabilistic key distillation in a more strict sense, where the distillation protocol must yield a perfect secret key whenever the protocol is successful, even if it requires an arbitrarily large number of resource states.
@@
One of the main tools that we use to establish our results is the resource theory of unextendibility [23, 24] and its state-dependent variation [25]. This resource theory was developed in [23, 24] as a relaxation of the resource theory of entanglement, in which one-way LOCC channels are allowed for free. An important quantity that we employ is the min-unextendible entanglement of a bipartite state [25]. Our work also provides significant improvements on existing bounds [25, Sec. V.D] regarding the overhead of probabilistic one-way secret key distillation.
2 One-way secret-key distillation
Let us begin by considering the task of secret-key distillation. The objective of such a protocol $\mathcal{L}_{AB\to A^{\prime}B^{\prime}}$ is to transform a bipartite state $\rho_{AB}$, purified by $\psi_{ABE}$, into a tripartite key state as follows:
@@ 2 One-way secret-key distillation
where $\omega_{A_{1}B_{1}}$ is an arbitrary quantum state and $\left(U_{i}\right)_{i=0}^{k-1}$ is a tuple of unitary operators. The systems $A_{0}B_{0}$ are the key systems, and $A_{1}B_{1}$ are the shield systems. The bipartite private state defined in (2) can be used to distill at least $\log_{2}k$ secret key bits. We can hence reframe the task of secret key distillation into the distillation of bipartite private states. From here on we simplify the labeling for the key systems and shield systems: when referring to a private state $\gamma^{k}_{A_{0}A_{1}B_{0}B_{1}}$, we group the systems held by Alice into a single system label $A\coloneqq A_{0}A_{1}$ and all the systems held by Bob into a single system label $B\coloneqq B_{0}B_{1}$.
@@
where we have used the shorthand
$i]\coloneqq|i\rangle\!\langle i|.$ (4)
In (3), $\gamma^{k}_{A^{\prime}B^{\prime}}$ is a bipartite private state with at least $\log_{2}k$ bits of secrecy, system $X$ is a classical flag indicating the success or failure of the protocol, and $\sigma_{A^{\prime}B^{\prime}}$ is an arbitrary quantum state generated when the protocol fails to distill a private state.
In an arbitrary probabilistic secret-key distillation protocol, both parties must have access to the flag $X$ in order to use the distilled key for private communication. Suppose that a probabilistic secret-key distillation protocol fails to establish a secret key. In that case, both parties involved in the distillation process can discard their systems and repeat the protocol with another instance of the resource state. If one-way LOCC channels are available for free, it suffices to demand that Alice receives the flag $X$ since she can send the flag to Bob using the freely available forward classical channel.
@@
In (6) we require that the state generated upon failure of the protocol is a maximally mixed state. This additional constraint on secret-key distillation protocols does not affect the maximum expected number of secret key bits that can be distilled from a bipartite state using one-way LOCC channels, as we show in Appendix A.
It is worthwhile to note that a maximally entangled state of Schmidt rank $k$ is a private state holding $\log_{2}k$ secret key bits [9, 10]. As such, an arbitrary entanglement distillation protocol can be transformed into a secret-key distillation protocol without affecting the rate of distillation. Hence, the one-way distillable secret key of a quantum state is not less than the one-way distillable entanglement of the state in both the probabilistic and approximate settings.
A simpler way to analyze the probabilistic distillation of secret keys under the action of one-way LOCC channels is by considering the erasure symbol. The erasure symbol $[e]$ is defined to be orthogonal to every state in the span of $\{|i\rangle\!\langle j|\}_{i,j=0}^{d-1}$, where $d$ is the dimension of the underlying system. For practical purposes, one can think of the erasure symbol as a pure state in the $(d+1)$-dimensional Hilbert space that is orthogonal to every state in the $d$-dimensional Hilbert space representing the system of interest. In the case of a joint system $S$ that comprises multiple subsystems, say, $S_{1},S_{2},\ldots,S_{k}$, the erasure symbol $e]_{S}$ can be represented as follows:
$e]_{S}=[e]_{S_{1}}\otimes[e]_{S_{2}}\otimes\cdots\otimes[e]_{S_{k}},$ (7)
which is orthogonal to every state on the system $S$, and hence, it is consistent with the definition of the erasure symbol.
In a one-way secret-key distillation protocol, if Alice finds the flag $X_{A}$ in the state $0]_{X_{A}}$, she can erase her state, and she can instruct Bob to erase his state as well. We call the resulting state the doubly erased private state, which has the following form:
$\eta^{p,k}_{AB}\coloneqq p\penalty 10000\ \gamma^{k}_{AB}+(1-p)[e]_{A}\otimes[e]_{B},$ (8)
where both Alice and Bob can retrieve the flag by performing the POVM $\{\Pi,[e]\}$ on their respective systems, where $\Pi\coloneqq\sum_{i=0}^{d-1}[i]$. They can further replace their state with a maximally mixed state upon measuring the erasure symbol, hence retrieving the distilled state in (6). That is,
$\eta^{p,k}_{AB}\xleftrightarrow[]{\operatorname{LO}}p[1]_{X_{A}}\otimes\gamma^{k}_{A^{\prime}B^{\prime}}+(1-p)[0]_{X_{A}}\otimes\frac{I_{A^{\prime}B^{\prime}}}{d_{A^{\prime}}d_{B^{\prime}}}.$ (9)
@@
3 Min-unextendible entanglement
The min-unextendible entanglement of a bipartite state was defined in [25, Sec. 4.2] as a monotone for the state-dependent resource theory of unextendibility. We briefly mention the relevant properties of this quantity here, and we refer the reader to [25] for a detailed presentation of the resource theory of unextendibility.
The min-unextendible entanglement is defined with respect to the min-relative entropy [26, Def. 2] as follows:
@@ 3 Min-unextendible entanglement
$\mathcal{F}\!\left(\rho_{AB}\right)\coloneqq\left\{\begin{array}[]{c}\operatorname{Tr}_{B}\!\left[\omega_{ABB^{\prime}}\right]:\par\rho_{AB}=\operatorname{Tr}_{B^{\prime}}\!\left[\omega_{ABB^{\prime}}\right],\\ \omega_{ABB^{\prime}}\in\mathcal{S}\!\left(ABB^{\prime}\right)\end{array}\right\},$ (11)
with $\mathcal{S}\!\left(ABB^{\prime}\right)$ the set of all states of the joint system $ABB^{\prime}$ and system $B^{\prime}$ being isomorphic to system $B$.
@@
$E^{u}_{\min}\!\left(\eta^{p,k}_{AB}\right)=\\ E^{u}_{\min}\!\left(p[1]_{X_{A}}\otimes\gamma^{k}_{AB}+(1-p)[0]_{X_{A}}\otimes\frac{I_{AB}}{d_{A}d_{B}}\right).$ (12)
4 Limitations on probabilistic one-way distillable secret key
The monotonicity of the min-unextendible entanglement of a bipartite state under the action of one-way LOCC channels implies that the min-unextendible entanglement of the target state in a probabilistic one-way secret-key distillation protocol does not exceed the min-unextendible entanglement of the source state. We first present a lower bound on the min-unextendible entanglement of the doubly erased private state, which is the target state of a probabilistic one-way secret-key distillation protocol.
@@ 4 Limitations on probabilistic one-way distillable secret key
The min-unextendible entanglement of the doubly erased private state $\eta^{p,k}_{AB}$ is strictly positive for all $p\in(0,1]$ and every integer $k\geq 2$, and it is only equal to zero at $p=0$ because the resulting state is a product state ([25, Proposition 3]). As a consequence of the one-way LOCC monotonicity of the min-unextendible entanglement, a state whose min-unextendible entanglement is equal to zero cannot be used to distill any number of secret key bits with a non-zero probability using one-way LOCC channels. In what follows, we identify the structure of the set of states for which the min-unextendible entanglement is equal to zero and state the main result of this work formally in Theorem 1.
Let us first analyze the set of states whose min-unextendible entanglement is equal to zero. If there exists a state $\sigma_{AB}\in\mathcal{F}\!\left(\rho_{AB}\right)$ such that $\rho_{AB}=\sigma_{AB}$, then $\rho_{AB}$ is a two-extendible state [27, 28, 29], and its min-unextendible entanglement is equal to zero. A more general set of states for which the min-unextendible entanglement is equal to zero, which we call super two-extendible states, can be defined as follows:
@@
The set of super two-extendible states is convex but not closed (see Appendix D).
@@
While the approximate one-way distillable secret key of a two-extendible state, also known as an anti-degradable state [30], is equal to zero [31, Thm. 15.43], the same is not true for a general super two-extendible state, as we shall see later in this paper. However, a super two-extendible state cannot be used for probabilistic one-way secret key distillation since its min-unextendible entanglement is equal to zero. Combining this fact with the additive property of the min-unextendible entanglement, we arrive at our main no-go theorem stated as Theorem 1, which identifies a broad set of states for which the probabilistic one-way distillable secret key is equal to zero.
@@
4.1 Erased private state
Let us consider the following state, which is established if Bob’s share of a bipartite private state $\gamma^{k}_{AB}$ gets erased with probability 1p1-p:
$\widetilde{\eta}^{p,k}_{AB}\coloneqq p\penalty 10000\ \gamma^{k}_{AB}+(1-p)\operatorname{Tr}_{B}\!\left[\gamma^{k}_{AB}\right]\otimes[e]_{B},$ (15)
@@ 4.1 Erased private state
$\widetilde{\Phi}^{p,d}_{AB}\coloneqq p\penalty 10000\ \Phi^{d}_{AB}+(1-p)\frac{I_{A}}{d_{A}}\otimes[e]_{B},$ (16)
@@
The set of full-rank states also lies inside the set of super two-extendible states, which leads to Corollary 2 below.
@@
The inability to probabilistically distill resources from full-rank states has been observed in general quantum resource theories [32, 33, 22]. Corollary 2 extends this result to probabilistic secret-key distillation under one-way LOCC channels. This result also implies that the overhead of probabilistic secret-key distillation considered in [25, Sec. V-B] is infinite for quantum states with full-rank density matrices, thus providing a significant strengthening of the bounds from [25, Sec. V-B].
4.2 Approximate versus probabilistic secret key distillation
Given the inability to distill a secret key from super two-extendible states even probabilistically, it is natural to ask if one can distill a key from such states if an arbitrarily small error is allowed. The rate at which secret-key bits can be distilled from a state such that the error vanishes asymptotically, using a one-way LOCC protocol, is called the one-way distillable key of the state [12], or the approximate one-way distillable key of the state, as we call it in this paper to clearly distinguish it from the probabilistic one-way distillable key of the state.
Indeed, it is possible distill a secret key from a state when an asymptotically vanishing error is allowed in distillation even if a perfect key cannot be distilled with any non-zero probability from the states, as we can see from the example of erased states. The approximate one-way distillable secret key is not less than the approximate one-way distillable entanglement of a state, which in turn is bounded from below by the coherent information of the state [12, Thm. 10]. For $p>\frac{1}{2}$, the coherent information of an erased state $\widetilde{\Phi}^{p,d}_{AB}$ is strictly non-negative. Similarly, the coherent information of some isotropic [34] and Werner states [35], which are full-rank states, is non-zero (see Appendices I and J). However, Corollaries 1 and 2 state that the probabilistic one-way distillable secret key of these states is equal to zero, demonstrating an extreme gap between the probabilistic and approximate one-way distillable secret key.
5 Conclusion
We found a convex set of states, that we call super two-extendible states, from which it is impossible to probabilistically distill perfect secret-key bits with any non-zero probability. Our main result is a no-go theorem stating that the probabilistic one-way distillable secret key of a super two-extendible state, and consequently its probabilistic one-way distillable entanglement, is equal to zero.
We demonstrated an extreme gap between the probabilistic and approximate one-way distillable secret key for some states. As such, we showed that probabilistic distillation of perfect secret keys using one-way LOCC protocols is not feasible for most states of interest, and one must allow some error for practically reasonable key distillation protocols. Considering the intermediate regime between probabilistic and approximate distillation [33, 21, 36], where the goal is to distill high-fidelity maximal-resource states with a non-zero probability, would shed light on the trade-off between distillation error and success probability in the task of probabilistic distillation of noisy secret keys.
Acknowledgments